Project

General

Profile

NDOS Commissioning Meeting 2011/02/21

Attendance: Jon, Andrew, Peter, Kurt, Paul, Mat, Alex, Minerba, Gavin, Susan, Alena, Rick K.

Agenda:

  1. Channel Recovery Plans
  2. DAQ Update & Plans
  3. DCS Status & Plans
  4. Shifters' report
  5. AOB

NDOS Building Schedule:

Channel recovery (Mat):

  • 33 APDs were removed on Friday, 10 new were installed onto side of block 2
  • 10 more were installed on Saturday on top of diblock 2
  • New pixel masks will be generated soon
  • 15 more heatsinks are available (20 APDs available). More heatsinks are on their way. These will likely be installed tomorrow (Mat + Minerba)

Plan for electronics:

  • try first voltage regulation
  • next try individual thresholds

DAQ Software:

  • Kurt loaded historical DAQ configuration data for runs 10738-11555
  • RC bug fixed Friday night
  • More RC bug fixes will go in later this week
  • DDS upgrade is not quite ready yet.
  • More DCM kernel and debugging tools also later this week.
  • DCM register readback (Peter):
  • tested last week, seemed to work
  • monitor program is being worked on by Jon, hopefully something will be ready later this week

DCS:

  • Jon requested more temperature readouts on different catwalk levels

Any Other Business: